OpenStudy (anonymous):

What is the equation of the circle with center (2,6) radius 5?

OpenStudy (anonymous):

the equation of a circle is: (x-h)^2+(y-k)^2 = r^2 where (h,k) is a point on the circle and r is the radius what can you plug into the formula with the given information?

OpenStudy (calculusfunctions):

@Maymay07 The equation of a circle with center (h, k) is given by\[(x -h)^{2}+(y -k)^{2}=r ^{2}\]where (h, k) is the center of the circle, NOT a point on the circle as @etemplin stated, and r is the radius of the circle.

OpenStudy (anonymous):

oh...sorry my mistake

OpenStudy (calculusfunctions):

That's OK, it happens but you must pay more attention to detail, especially when you're trying to teach someone.

OpenStudy (anonymous):

Thank you :)